A person drops a cylindrical steel bar (Y = 1.000
Answer #1

Apply conservation of energy between top an bottom positions

mg(h+x) = 0.5*k*x^2

from the Young modulus relation

K = YA /L

= 1 * 10^11 * 3.14*0.0065^2 / 0.72

= 1.842 x 10^7 N / m

---------------------------------------

0.6(9.8)(1.5+x) = 0.5*1.842 x 10^7 N / m*x^2

9210000 x^2 -5.88 x -8.82 = 0

by solving quadratic equation

x = 9.78 * 10^-4 m